# RingMoE: Mixture-of-Modality-Experts Multi-Modal Foundation Models for Universal Remote Sensing Image Interpretation

### v1

@@ B.4.3 Multi-modal Semantic Segmentation
  • WHU-OPT-SAR (Opt/SAR-L2) [74], is a multi-modal scene-level semantic segmentation dataset includes 100 pairs of Opt and amplitude SAR images, with each pair representing the same geographic area. The images have a spatial resolution of 5 meters and dimensions of 5556$\times$3704. The dataset features seven categories: “Farmland”, “City”, “Village”, “Water”, “Forest”, “Road”, and “Others”. Following the official dataset guidelines, the images are cropped into 256$\times$256 pixel patches without overlapping. The dataset is split into training and testing sets with an 8:2 ratio.
  • DFC23 (Opt/SAR-L2) [75], is a large-scale, multi-modal dataset designed for building roof type classification and includes both Opt and amplitude SAR images. Captured by the SuperView-1, Gaofen-2, and Gaofen-3 satellites with spatial resolutions of 0.5, 0.8, and 1 meter, respectively, the dataset spans 12 roof geometry categories, such as “Flat roof”, “Gable roof”, “Gambrel roof”, “Row roof”, and multiple roof types. For experiments, images are cropped to 512$\times$512 pixels, and the dataset is split into training, validation, and testing sets with a 7:2:1 ratio.
(2) Implementation Detail. The multi-modal segmentation downstream tasks are implemented using the mmsegmentation framework. We employ Upernet [86] as the segmentation network, with RingMoE-Opt and RingMoE-SARL2 serving as feature extractors for optical and amplitude SAR images, respectively. After extracting features from both modalities, we concatenate them and apply 3$\times$3 and 1$\times$1 convolutional layers for straightforward feature fusion, leveraging the inherent representational strengths of RingMoE. Finally, segmentation predictions are generated using the UperNet Head.
The optimizer is AdamW with an initial learning rate of 6e-5 and a weight decay of 0.01. The learning rate schedule consists of two phases: a linear warm-up for the first 1,500 iterations with a factor of 0.1, followed by a polynomial scheduler with a minimum learning rate of 1e-6. For the WHU-OPT-SAR dataset, the images are cropped to 256$\times$256, with training set at 80k iterations and a batch size of 16. For the DFC23 dataset, the images are cropped to 512$\times$512, with training set at 80k iterations and a batch size of 12. In our experiments, we employ regular data augmentation strategies, including RandomFlip, RandomScaling (from 0.5 to 2.0), and RandomCrop.
@@ B.4.4 Object Detection
  • DIOR [77] and DIOR-R (Opt) [87], is a large-scale benchmark for object detection in optical remote sensing. It comprises 23,463 images of 800$\times$800 pixels, sourced from over 80 countries. The dataset includes annotations for 192,472 instances across 20 object classes, each labeled with horizontal bounding boxes. The dataset is divided into 5862 images for training, 5863 images for validation, and 11738 images for testing. Following [16, 27], we utilize the training and validation sets together for training, and the test set for evaluation. DIOR-R is an enhanced version of DIOR, containing the same images but with oriented bounding box annotations, adding complexity to the detection task. Similar to DIOR, we merge the training and validation sets for training and reserve the test set for performance evaluation.
  • HRSC2016 (Opt) [78], is a public oriented ship detection dataset comprising images from Google Earth of diverse sizes, spanning from 300$\times$300 to 1500$\times$900, and most of them are about 1000$\times$600. The dataset includes 2976 objects with oriented bounding box annotations extracted from six important ports. The images are divided into 436 for training, 181 for validation, and 444 for testing.
@@ B.4.5 Object Tracking
  • AIR-MOT (Opt) [81], is an object tracking dataset consisting of 152 videos captured by the Jilin-1 satellite between October 2017 and October 2020. The videos have a resolution of 1920$\times$1080 pixels and a frame rate ranging from 5 to 10 FPS, with each video containing over 70 timestamps. The dataset includes 5736 instances of two object categories (“Aircraft” and “Ship”), collected from ten diverse regions across the globe. The complex and varied backgrounds, along with the multiscale characteristics of objects, make it challenging for trackers to maintain robustness. The dataset is split into 106 training videos and 46 testing videos. We evaluate model performance on the test set.
  • AIR-HSAO (Opt) [82], is a high-speed aerial object tracking dataset built from satellite videos captured by the Jilin satellite between 2016 and 2022. The dataset consists of 197 video clips, with a resolution of 1920$\times$1080 pixels, and includes 1437 airplane trajectories. The dataset focuses on fast-moving airplanes, with video clips covering various regions, such as urban areas, airports, and harbors. Object sizes range from less than 200 to over 5500 pixels, showcasing multiscale characteristics. The diverse and complex backgrounds add to the challenge of object tracking. The dataset is split into 128 training videos and 69 testing videos. We evaluate model performance on the test set.
(2) Implementation Detail. All object tracking tasks utilize the mmdetection framework, with ByteTrack [^5] https://github.com/ifzhang/ByteTrack as the tracking algorithm. The training process follows a consistent strategy across two datasets. First, a two-stage training approach is applied to the detection model, using compressed RingMoE-Opt (KC) as the backbone for feature extraction and standard YOLOX as the detection framework. In the first stage, video frames are cropped to 640$\times$640 pixels for comprehensive fine-tuning of all parameters. The AdamW optimizer is employed with an initial learning rate of 1e-5, a weight decay of 0.005, and a batch size of 2. The learning rate linearly warms up over the first 5,000 iterations by a factor of 0.1, then remains fixed for a total of 5 epochs. In the second stage, the backbone parameters are frozen, and fine-tuning is performed on the original 1920$\times$1080 resolution frames, using the SGD optimizer with a learning rate of 4e-4, a weight decay of 0.005, and a batch size of 2, for an additional 2 epochs. For tracking inference, ByteTrack operates on the trained detection model with the following settings: a non-maximum suppression (NMS) threshold of 0.65, a tracking threshold of 0.3, a track buffer of 50 frames for maintaining lost tracks, and a matching threshold of 0.8.
@@ B.4.6 Change Detection
  • CDD (Opt) [84], is a bitemporal general change detection benchmark focused on capturing seasonal variations in remote sensing images. It consists of 16,000 image pairs, each 256$\times$256 pixels, with spatial resolutions ranging from 0.03 to 1 meter per pixel. The dataset is split into 10,000 pairs for training and 3,000 pairs for testing and validation, offering a comprehensive foundation for evaluating change detection techniques under varying spatial and seasonal conditions.
(2) Implementation Detail. For all change detection tasks, we employ the Bidirectional Integration Transformer (BIT) [^6] https://github.com/justchenhao/BIT [90] framework, integrating the compressed RingMoE-Opt (KC) model as the backbone for feature extraction. Training is optimized with the AdamW optimizer and binary cross-entropy loss. For efficient fine-tuning, the parameters of the first stage in RingMoE are frozen, while the remaining three stages undergo training. The initial learning rate is set to 5e-5 and decays gradually to zero over 200 epochs. Data augmentation techniques include flipping, cropping, re-scaling, color jittering, and Gaussian blur. Given computational constraints, all images are uniformly divided into non-overlapping slices of 256$\times$256 pixels for training, validation, and testing.
[…]
B.4.7 Depth Estimation
@@
  • ISPRS Vaihingen, is a widely used dataset for semantic segmentation and depth estimation tasks. The dataset contains 33 high-resolution images with a spatial resolution of 0.09 meters. In terms of data composition, the dataset contains three-channel (near-infrared, red, and green) images and corresponding digital surface model (DSM) data. Following the original experimental settings, we divide it into 16 training images and 17 test images during training and testing, and the digital surface model corresponding to each image is used as the label data for training.
  • ISPRS Potsdam, is similar to the Vaihingen dataset, which consists of 38 high-resolution images with a spatial resolution of 0.05 meters. The dataset contains four-channel (near-infrared, infrared, green, and blue) high-resolution images and corresponding digital surface model (DSM) data. Similarly, following the original dataset setting, we divide the Potsdam dataset into 24 training images and 14 testing images during training and testing. In addition, for the convenience of training, we did not consider the near-infrared channel and only input red, green, and blue as channels into the network.
(2) Implementation Detail. The experiments employ the Binsformer [^7] https://github.com/zhyever/Monocular-Depth-Estimation-Toolbox [91] algorithm, utilizing RingMoE-Opt (KC) as the backbone network. Images are cropped to size 512$\times$512 for input, and data augmentation is performed using random rotation and flipping. The model is trained with a batch size of 8 and an initial learning rate of 1e-4 for 38.4k iterations. The AdamW optimizer is used with parameters ($\beta_{1}$, $\beta_{2}$, wd) = (0.9, 0.999, 0.01), where wd denotes the weight decay. A linear learning rate warm-up strategy is applied over the first 30% of the total iterations to stabilize early training.
@@
{tblr}
cells = c, cell21 = r=3, cell51 = r=3, hline1-2,5,8 = -, hline3-4,6-7 = 2-4, Datasets & Fold-i Training classes Testing classes
@@
B.4.8 Unseen Class Recognition
@@
  • NWPU-RESISC45 (classification) [60], is a publicly available dataset for RS scene classification. It contains 31,500 RGB images with a resolution of 256$\times$256 pixels. The dataset includes a total of 45 scene categories, with each category comprising 700 images. Following the settings in [92, 93], the dataset is divided into 25, 10, and 10 categories for training, validation, and testing, respectively, as detailed in Tab.A1.
  • WHU-RS19 (classification) [61], is released by Wuhan University in 2012. It comprises 1,005 RGB images across 19 scene categories, with a minimum of 50 images per category. Each image is 600$\times$600 pixels in size. Following the settings in [92, 93], the dataset is divided into 9, 5, and 5 categories for training, validation, and testing, respectively, as detailed in Tab.A1.
  • UCMerced (classification) [62], is designed for land-use RS scene recognition and includes 21 distinct scene categories, such as “Agriculture”, “Airplane”, “Baseball field”, and “Beach”. Each category contains 100 images with a resolution of 256$\times$256 pixels. Following the settings in [92, 93], the dataset is divided into 10, 6, and 5 categories for training, validation, and testing, respectively, as detailed in Tab.A1.
  • iSAID (segmentation) [63], is a large-scale benchmark for evaluating instance segmentation and semantic segmentation algorithms, containing 655,451 object instances from 2,806 high-resolution images. It includes 15 object categories, such as “Ship”, “Baseball diamond”, and “Airplane”. To assess the model’s generalization capability on unseen classes, following the setup in [94, 95], the 15 categories are divided into three groups (Fold-i), each consisting of 10 training classes and 5 testing classes, ensuring no overlap between training and testing classes within the same group. This cross-division approach leverages the dataset’s diversity to thoroughly evaluate the model’s generalization performance across different categories. Detailed class divisions are provided in Tab.A2.
@@
Specialized method RelationNet(CVPR’2018) [58] 66.21±0.28 61.74±0.51 48.48±0.75
@@
Foundation model RingMo(TGRS’2023) [16] 72.88±0.23 74.47±0.37 57.67±0.74
@@
TABLE A4: Detailed performance comparison of RingMoE with varying parameter configurations on Few-shot Segmentation tasks, i.e., iSAID and LoveDA. Bold and underline denote the best and second-best results, respectively.
@@
Specialized method PFENet(TPAMI’2022) [104] 47.30 22.10
@@
Foundation model RingMoE(6.5B) 52.45 32.40
@@
(2) Implementation Detail. For the few-shot classification task, we conduct evaluations under the 5-way 1-shot and 5-way 5-shot settings. Specifically, the task requires classifying the query images into one of five novel classes, where each novel class is guided by either one labeled (1-shot) or five labeled examples (5-shot). Following the RelationNet [58], we leverage RingMoE-KC (frozen) for feature extraction, obtaining features for both the query images and the few labeled examples from the novel classes. Then we utilize the similarity metric to calculate the score between the query image and the labeled samples for classification. The input images from both datasets are resized to 256$\times$256 pixels and underwent data augmentation, including random cropping and horizontal flipping. The SGD optimizer is used with a learning rate of 0.001 and a weight decay of 5e-4.
For the few-shot segmentation task, we evaluate under the 1-shot setting, where a labeled sample containing a novel class (namely support image) is provided to guide the segmentation of the novel class objects in the query image. We utilize RingMoE-KC to extract features from both the support and query images. Then, we apply the foreground mask of the support image to perform the masked average pooling operation, obtaining both the foreground and background prototypes. Finally, we calculate the similarity between the query features and the support prototypes to achieve segmentation. For iSAID, the image size is resized to 256$\times$256, while for LoveDA, the image size is resized to 384$\times$384. We employ SGD for optimization with a weight decay of 0.0001 and a momentum of 0.9. The batch size is set to 8, and the model is trained for 60 epochs. The initial learning rate for iSAID is 0.005, while for LoveDA, it is 0.003.
@@ C.2 Performance on Scene Classification
Tab.2 in the main body presents the performance comparison of the proposed RingMoE and other foundation models on two widely used RS scene classification datasets. The results demonstrate that RingMoE achieves outstanding performance even with limited data. Specifically, the 6.5B-parameter RingMoE achieves 95.90% and 98.19% accuracies on NWPU-RESISC45 and AID, respectively, outperforming SkySense [27] by 1.05% and 0.51%, and Scale-MAE [17] by 3.27% and 1.75%.

### v2

@@ B.5.3 Multi-modal Semantic Segmentation
[…]
  • DFC23 (Opt/SAR-L2) [76], is a large-scale, multi-modal dataset designed for building roof type classification and includes both Opt and amplitude SAR images. Captured by the SuperView-1, Gaofen-2, and Gaofen-3 satellites with spatial resolutions of 0.5, 0.8, and 1 meter, respectively, the dataset spans 12 roof geometry categories, such as “Flat roof”, “Gable roof”, “Gambrel roof”, “Row roof”, and multiple roof types. For experiments, images are cropped to $\times512$ pixels, and the dataset is split into training, validation, and testing sets with a 7:2:1 ratio.
(2) Implementation Detail. The multi-modal segmentation downstream tasks are implemented using the mmsegmentation framework. We employ Upernet [93] as the segmentation network, with RingMoE-Opt and RingMoE-SARL2 serving as feature extractors for optical and amplitude SAR images, respectively. After extracting features from both modalities, we concatenate them and apply $\times3$ and $\times1$ convolutional layers for straightforward feature fusion, leveraging the inherent representational strengths of RingMoE. Finally, segmentation predictions are generated using the UperNet Head.
The optimizer is AdamW with an initial learning rate of 6e-5 and a weight decay of 0.01. The learning rate schedule consists of two phases: a linear warm-up for the first 1,500 iterations with a factor of 0.1, followed by a polynomial scheduler with a minimum learning rate of 1e-6. For the WHU-OPT-SAR dataset, the images are cropped to $\times256$, with training set at 80k iterations and a batch size of 16. For the DFC23 dataset, the images are cropped to $\times512$, with training set at 80k iterations and a batch size of 12. In our experiments, we employ regular data augmentation strategies, including RandomFlip, RandomScaling (from 0.5 to 2.0), and RandomCrop.
@@ B.5.4 Object Detection
  • DIOR [78] and DIOR-R (Opt) [94], is a large-scale benchmark for object detection in optical remote sensing. It comprises 23,463 images of $\times800$ pixels, sourced from over 80 countries. The dataset includes annotations for 192,472 instances across 20 object classes, each labeled with horizontal bounding boxes. The dataset is divided into 5862 images for training, 5863 images for validation, and 11738 images for testing. Following [15, 92], we utilize the training and validation sets together for training, and the test set for evaluation. DIOR-R is an enhanced version of DIOR, containing the same images but with oriented bounding box annotations, adding complexity to the detection task. Similar to DIOR, we merge the training and validation sets for training and reserve the test set for performance evaluation.
[…]
@@ B.5.5 Object Tracking
[…]
@@ B.5.6 Change Detection
  • CDD (Opt) [85], is a bitemporal general change detection benchmark focused on capturing seasonal variations in remote sensing images. It consists of 16,000 image pairs, each $\times256$ pixels, with spatial resolutions ranging from 0.03 to 1 meter per pixel. The dataset is split into 10,000 pairs for training and 3,000 pairs for testing and validation, offering a comprehensive foundation for evaluating change detection techniques under varying spatial and seasonal conditions.
(2) Implementation Detail. For all change detection tasks, we employ the Bidirectional Integration Transformer (BIT) [^6] https://github.com/justchenhao/BIT [97] framework, integrating the compressed RingMoE-Opt (KC) model as the backbone for feature extraction. Training is optimized with the AdamW optimizer and binary cross-entropy loss. For efficient fine-tuning, the parameters of the first stage in RingMoE are frozen, while the remaining three stages undergo training. The initial learning rate is set to 5e-5 and decays gradually to zero over 200 epochs. Data augmentation techniques include flipping, cropping, re-scaling, color jittering, and Gaussian blur. Given computational constraints, all images are uniformly divided into non-overlapping slices of $\times256$ pixels for training, validation, and testing.
(3) Evaluation Metric. We employ the F1 score to evaluate the accuracy of identifying changed areas between two images. It is the harmonic mean of precision and recall, where precision measures the proportion of correctly detected changes out of all detected changes, and recall measures the proportion of correctly detected changes out of all actual changes. It can be formulated as $\text{F1}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$, where $\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ and $\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$.
B.5.7 Depth Estimation
@@
  • ISPRS Vaihingen [71], is a widely used dataset for semantic segmentation and depth estimation tasks. The dataset contains 33 high-resolution images with a spatial resolution of 0.09 meters. In terms of data composition, the dataset contains three-channel (near-infrared, red, and green) images and corresponding digital surface model (DSM) data. Following the original experimental settings, we divide it into 16 training images and 17 test images during training and testing, and the digital surface model corresponding to each image is used as the label data for training.
  • ISPRS Potsdam [71], is similar to the Vaihingen dataset, which consists of 38 high-resolution images with a spatial resolution of 0.05 meters. The dataset contains four-channel (near-infrared, infrared, green, and blue) high-resolution images and corresponding digital surface model (DSM) data. Similarly, following the original dataset setting, we divide the Potsdam dataset into 24 training images and 14 testing images during training and testing. In addition, for the convenience of training, we did not consider the near-infrared channel and only input red, green, and blue as channels into the network.
(2) Implementation Detail. The experiments employ the Binsformer [^7] https://github.com/zhyever/Monocular-Depth-Estimation-Toolbox [98] algorithm, utilizing RingMoE-Opt (KC) as the backbone network. Images are cropped to size $\times512$ for input, and data augmentation is performed using random rotation and flipping. The model is trained with a batch size of 8 and an initial learning rate of 1e-4 for 38.4k iterations. The AdamW optimizer is used with parameters ($\beta_{1}$, $\beta_{2}$, wd) = (0.9, 0.999, 0.01), where wd denotes the weight decay. A linear learning rate warm-up strategy is applied over the first 30% of the total iterations to stabilize early training.
@@
Datasets Fold-i Training classes Testing classes
@@
B.5.8 Unseen Class Recognition
@@
  • NWPU-RESISC45 (classification) [58], is a publicly available dataset for RS scene classification. It contains 31,500 RGB images with a resolution of $\times256$ pixels. The dataset includes a total of 45 scene categories, with each category comprising 700 images. Following the settings in [99, 100], the dataset is divided into 25, 10, and 10 categories for training, validation, and testing, respectively, as detailed in Tab.A2.
  • WHU-RS19 (classification) [59], is released by Wuhan University in 2012. It comprises 1,005 RGB images across 19 scene categories, with a minimum of 50 images per category. Each image is $\times600$ pixels in size. Following the settings in [99, 100], the dataset is divided into 9, 5, and 5 categories for training, validation, and testing, respectively, as detailed in Tab.A2.
  • UCMerced (classification) [60], is designed for land-use RS scene recognition and includes 21 distinct scene categories, such as “Agriculture”, “Airplane”, “Baseball field”, and “Beach”. Each category contains 100 images with a resolution of $\times256$ pixels. Following the settings in [99, 100], the dataset is divided into 10, 6, and 5 categories for training, validation, and testing, respectively, as detailed in Tab.A2.
  • iSAID (segmentation) [61], is a large-scale benchmark for evaluating instance segmentation and semantic segmentation algorithms, containing 655,451 object instances from 2,806 high-resolution images. It includes 15 object categories, such as “Ship”, “Baseball diamond”, and “Airplane”. To assess the model’s generalization capability on unseen classes, following the setup in [101], the 15 categories are divided into three groups (Fold-i), each consisting of 10 training classes and 5 testing classes, ensuring no overlap between training and testing classes within the same group. This cross-division approach leverages the dataset’s diversity to thoroughly evaluate the model’s generalization performance across different categories. Detailed class divisions are provided in Tab.A3.
@@
Specialized
method
RelationNet(CVPR’2018) [56] 66.21±0.28 61.74±0.51 48.48±0.75
@@
Foundation
model
RingMo(TGRS’2023) [15] 72.88±0.23 74.47±0.37 57.67±0.74
@@
TABLE A5: Detailed performance comparison of RingMoE with varying parameter configurations on Few-shot Segmentation tasks, i.e., iSAID and LoveDA. Note that SegGPT [109] was pre-trained with access to all categories from both datasets, while the other methods maintain evaluations on unseen classes. Bold and underline denote the best and second-best results, respectively.
@@
Specialized
method
PFENet(TPAMI’2022) [110] 47.30 22.10
@@
Foundation
model
SAM(ICCV’2023) [63] 45.38 17.11
SegGPT(ICCV’2023) [109] 52.16 34.27
VRP-SAM(CVPR’2024) [64] 49.03 19.32
RingMoE(6.5B) 52.45 32.40
@@
(2) Implementation Detail. For the few-shot classification task, we conduct evaluations under the 5-way 1-shot and 5-way 5-shot settings. Specifically, the task requires classifying the query images into one of five novel classes, where each novel class is guided by either one labeled (1-shot) or five labeled examples (5-shot). Following the RelationNet [56], we leverage RingMoE-KC (frozen) for feature extraction, obtaining features for both the query images and the few labeled examples from the novel classes. Then we utilize the similarity metric to calculate the score between the query image and the labeled samples for classification. The input images from both datasets are resized to $\times256$ pixels and underwent data augmentation, including random cropping and horizontal flipping. The SGD optimizer is used with a learning rate of 0.001 and a weight decay of 5e-4.
For the few-shot segmentation task, we evaluate under the 1-shot setting, where a labeled sample containing a novel class (namely support image) is provided to guide the segmentation of the novel class objects in the query image. We utilize RingMoE-KC to extract features from both the support and query images. Then, we apply the foreground mask of the support image to perform the masked average pooling operation, obtaining both the foreground and background prototypes. Finally, we calculate the similarity between the query features and the support prototypes to achieve segmentation. For iSAID, the image size is resized to $\times256$, while for LoveDA, the image size is resized to $\times384$. We employ SGD for optimization with a weight decay of 0.0001 and a momentum of 0.9. The batch size is set to 8, and the model is trained for 60 epochs. The initial learning rate for iSAID is 0.005, while for LoveDA, it is 0.003.
@@ C.1 Performance on Unseen Class Recognition
Figure A4: Feature visualization comparison across RSFMs on Scene Classification tasks. We extract features from various RSFMs on the test sets of both classification datasets and employ t-SNE to visualize the feature distributions of 10 randomly selected categories, highlighting the feature clustering across different models.
@@
In contrast, general models like SAM [63] and SegGPT [109], although pretrained on massive RGB segmentation corpora, exhibit limited generalization to RS scenarios. SAM struggles with modality shifts, achieving only 45.38% and 17.11% mIoU, while SegGPT, despite being pre-exposed to all categories from both benchmarks, only marginally surpasses RingMoE on LoveDA (34.27%) but underperforms it on iSAID (52.16% vs. 52.45%). These results underscore the advantages of RingMoE’s modality-aware architecture and targeted pretraining for RS applications.
@@ C.2 Performance on Scene Classification
Tab.2 in the main body presents the performance comparison of the proposed RingMoE and other foundation models on two widely used RS scene classification datasets. The results demonstrate that RingMoE achieves outstanding performance even with limited data. Specifically, the 6.5B-parameter RingMoE achieves 95.90% and 98.19% accuracies on NWPU-RESISC45 and AID, respectively, outperforming SkySense [92] by 1.05% and 0.51%, and Scale-MAE [16] by 3.27% and 1.75%. In addition, we extract the features from various foundation models on the test set of the two datasets and randomly select 10 categories for visual representation using t-SNE (see Fig.A4). Results indicate that RingMoE exhibits a more discriminative representation capability than other foundation models, characterized by tighter intra-class clustering and larger inter-class variance.